OpenStudy (anonymous):

Will someone please help me solve? Algebra 2, multiple choice :( I really need some help...

OpenStudy (gorv):

yep

OpenStudy (gorv):

q plzzz

OpenStudy (anonymous):

Solve. \[(\sqrt{7})^{6x}=49^{x-6}\]

OpenStudy (anonymous):

\[A. x=-\frac{ 21 }{ 2 }\]

OpenStudy (anonymous):

B. x=-6

OpenStudy (anonymous):

\[C. x=-\frac{ 6 }{ 5}\]

OpenStudy (gorv):

wat is the power of sqrt 7

OpenStudy (anonymous):

D. x=-12

OpenStudy (anonymous):

6x?

OpenStudy (gorv):

D is the ans then

OpenStudy (gorv):

got it????

OpenStudy (anonymous):

Yep :)

OpenStudy (anonymous):

Thank you

OpenStudy (gorv):

my pleasure